OpenStudy (anonymous):

solve by elimination x-y+z=-1 x+y+3z=-3 2x-y+2z=0

OpenStudy (anonymous):

are you sdoing matrices?

OpenStudy (anonymous):

no its umm systems

OpenStudy (anonymous):

alright so with systems you know the 3 rules you can do right?

OpenStudy (anonymous):

elimination subsitution

OpenStudy (anonymous):

what you want to do in this case is take equation 1 and 3 and try to eliminate a variable by adding them together

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

and then take 1 and 2 and try to get them to cancel the same variable

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

soo 3x-2y+3z=-1

OpenStudy (anonymous):

not quite,alright so you have equation 1 and 2 x-y+z=-1 x+y+3z=-3 when dealing with systems you can add two equations and replace it with the row equation you are adding them with so if you were to add equation one with 2 you'd get x-y+z=-1 2x+4z=-4 2x-y+2z=0

OpenStudy (anonymous):

see how it cancels a variable y out?

OpenStudy (anonymous):

oh yea ok

OpenStudy (anonymous):

so what would why equal though?

OpenStudy (anonymous):

once you cancel the y variable in another equation you'll be able to get one variable and then find the others by using the equations you started with

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

your answers should be (x,y,z)=(4,2,-3)

OpenStudy (anonymous):

do you understand... i might be able to post a screen shot of my work if you want?